OpenStudy (anonymous):

Calc: Integrals with Trig integral (-pi)/(4) to (3pi)/(4) 8sec(theta)tan(theta) ? I know I first need to find the anti-derv but can I rewrite sec(theta) and tan(theta) using identities?

OpenStudy (jhannybean):

\[\large \int\limits_{ -\frac{\pi}{4}}^{\frac{3\pi}{4}} 8\sec(\theta)\tan(\theta)d(\theta)\] let \(\ u = (\theta)\) and factor out the 8.

OpenStudy (jhannybean):

\[\large 8\int\limits_{ -\frac{\pi}{4}}^{\frac{3\pi}{4}} \sec(u)\tan(u)d(u)\]

OpenStudy (anonymous):

Or...you can remember that \(\dfrac{d}{d\theta}\sec\theta=\sec\theta\tan\theta \implies \displaystyle \int\sec\theta\tan\theta \,d\theta= \sec\theta + C\). There really isn't a need to change the variable of integration.

OpenStudy (jhannybean):

Then you've got your identity \[\ \int sec(u)tan(u)du = sec(u)\] So now just integrate \[ \sec(\theta) ]_{-\frac{\pi}{4}}^{\frac{3\pi}{4}}\]

OpenStudy (jhannybean):

Well, I wasn't changing the variable, but just using it to fit the identity. :P

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

how do I Factor out the 8 ?

OpenStudy (jhannybean):

You can pull any constants out of an integral.

OpenStudy (jhannybean):

As long as they're multiplied to one function throughout. in this case they act as a scaling factor to sin(theta)cos(theta) therefore do not need to be integrated along with the function. Therefore they can be dealt with later once you integrate your function.

OpenStudy (anonymous):

ok so what exactly would be the anti-derivative in this case

OpenStudy (anonymous):

\[8\sec (\Theta)+C\]

OpenStudy (anonymous):

@ArkGoLucky, and then I can plug in the values on the integral? in this case, -pi/4 and 3pi/4? But shouldn't i change the sec(theta) to 1/cos(theta) ?

OpenStudy (anonymous):

@Jhannybean, thank you :)

OpenStudy (jhannybean):

Yw :) and no you don't need to change it to 1/cos(\(\theta\)) since you've already integrated your function :)

OpenStudy (anonymous):

ok thank you !